OpenStudy (anonymous):

Which expression is equivalent to a^m ÷ a^n? a^m − ^n a^m + ^n a^m ⋅ ^n a^m ÷^ n

OpenStudy (anonymous):

@texaschic101

OpenStudy (anonymous):

@undeadknight26

OpenStudy (anonymous):

@Hayleymeyer

OpenStudy (anonymous):

@Destinyyyy

Nnesha (nnesha):

when we divide same bases we should `subtract` their exponents

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so a

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

so a

Nnesha (nnesha):

yes right \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\]

OpenStudy (anonymous):

yes right

Nnesha (nnesha):

when you multiply same bases then you should `add` their exponents \[\huge\rm x^m \times x^n= x^{m +n}\]